OpenStudy (anonymous):

A pair of boots costs $50 and a certain coat costs $150. If the price of the boots is decreased by 20% and the price of the coat is decreased by 30%, which of the following is the sum of their new costs? $165 $200 $225 $265 $330

OpenStudy (ghazi):

20% of 50= 20/100*50= 0.2*50= 10 new cost of it = 50-10=$40 similarly 30% of 150 = 0.3*150=45 new cost of coat= 150-45=$105 and total new sum=105+40=$145

OpenStudy (anonymous):

that's not 1 of the choices, unfortunately :(

OpenStudy (ghazi):

i guess choices are wrong then...let me check this again

OpenStudy (anonymous):

i got the choices straight from the paper

OpenStudy (ghazi):

20% of 50= 0.2*50= 10 so new cost is 50-10=40 similarly 30% of 150= 0.3*150= 45 new cost is = 150-45=105 total new cost = 105+40=$ 145

OpenStudy (anonymous):

the lowest choice is $165

OpenStudy (anonymous):

$165 $200 $225 $265 $330

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

srry to bother u, bro :(

OpenStudy (ghazi):

it's cool :) ..i don't know..where i am making mistake....

OpenStudy (anonymous):

thanx anyways!

OpenStudy (ghazi):

:) YW

OpenStudy (anonymous):

@ghazi answer is correct.\[50*\frac{80}{100}+150*\frac{70}{100}=145 \]None of the answer alternatives are correct.

OpenStudy (ghazi):

@robtobey thanks....that's why i did it again...

OpenStudy (anonymous):

those were the only choices they gave me. srry

OpenStudy (ghazi):

it's cool there must be some printing error i guess 165 has been written at 145

OpenStudy (mathmate):

"If the answers are right, which OP insists, then the question may not be. Example: A pair of boots costs $50 and a certain coat costs $150. If the price of the boots is decreased by 10% and the price of the coat is decreased by 20%, which of the following is the sum of their new costs?" would give $165. Can you check the question? @mitchelsewbaran ?